A body moves from a position r²=(I + 2j –3k) meter under the influence of constant force F=( 3i +3–k) Newton. Calculate the work done by the force.




1
Expert's answer
2022-04-01T08:41:26-0400

Answer

Work done is dot product of force and position

"W=F.r\\=( 3i +3j–k) .(i+ 2j –3k) \\=3+6+3\\=12J"